OpenStudy (anonymous):

find the slope-intercept form of the line whose slope is 4 and that passes through the point (-3,6) The equation of the line is?

OpenStudy (anonymous):

y = 4x + b. Find b. Write that this line passes through point (-3, 6): 6 = 4(-3) + b -> b = .....

OpenStudy (anonymous):

can anyone help me please?

OpenStudy (mathmale):

@thu1935: thank you for contributing. However, the equation you posted is the slope-intercept form, NOT the point-slope form. @jamone: Hello! It'd be helpful if you would write down both the point-slope form and the slope-intercept form of the equation of a straight line. You are given the slope, m = 4, along with one point thru which your line passes, (-3,6). Substitute the coordinates of this point, and that slope, into \[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

(this latter equation is the "point-slope" form of the equation of a str. line.

OpenStudy (anonymous):

4x-y=18

OpenStudy (mathmale):

That may be right, Jamone, but please, show me the actual work you did in obtaining that answer. You didn't do what I asked you to do, which was to substitute the coordinates of the given point, along with the given slope, into the point-slope formula for a straight line.

OpenStudy (anonymous):

y-y1=m(x-x1) y-6=4(x-(-3) y-6=4x+3 y-6=4(x+3) y-6=4x+12 4x-y+6=12 4x-y=18

OpenStudy (mathmale):

thank you very much. Nice work. It's really worth going through all these steps. I think you've learned something valuable. Good going!

OpenStudy (mathmale):

I need to log off now, but hope to work with you again another time.

OpenStudy (anonymous):

thank you, have a good one.